OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0 Can someone help me please?

OpenStudy (anonymous):

to find x value?

OpenStudy (anonymous):

i think so

OpenStudy (precal):

why not just factor out sinx

OpenStudy (anonymous):

i just realized i typed that wrong, it's sin^2x+sinx=0

OpenStudy (precal):

\[\sin^2x+sinx=0\] \[sinx(sinx+1)=0\]

OpenStudy (precal):

\[sinx=0 \]

OpenStudy (anonymous):

so sinx=0 and-1?

OpenStudy (precal):

\[sinx+1=0\]

OpenStudy (precal):

\[sinx=-1\]

OpenStudy (precal):

did they give a restriction? example 0 to 2pi

OpenStudy (anonymous):

no, there is no restriction

OpenStudy (precal):

now look for where sinx=0 and sinx=-1

OpenStudy (precal):

then you need all solutions

OpenStudy (anonymous):

\[x=3/2\pi \]

OpenStudy (anonymous):

x= pi, x=3/2pi

OpenStudy (precal):

+2pik so that it has all of the solutions

OpenStudy (precal):

recall the period for sinx is 2pi so 2(pi)k will take care of all of the possibilities

OpenStudy (anonymous):

so would it be x=pi + 2(pi)k, x= 3/2pi + 2(pi)k

OpenStudy (precal):

I believe so, I think you can also include - so + and -

OpenStudy (anonymous):

ok thank you!!

OpenStudy (precal):

ok according to wolframalpha leave off the - just 2 pi (k)

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (precal):

yw